# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 4 Quantum algorithms
In this section, we introduce our quantum algorithms based on the LARS algorithm. In Section 4.1, we propose our simple quantum LARS algorithm that exactly computes the pathwise Lasso, while in Section 4.2, we improve upon this simple algorithm and propose the approximate quantum LARS algorithm. In Section 4.3, we dequantise the approximate LARS algorithm using sampling techniques. Finally, in Section 4.4, we analyse the algorithms’ complexity for the case when $\mathbf{X}$ is a standard Gaussian random matrix.
@@ 4.1 Simple quantum LARS algorithm
As a warm-up, we propose a simple quantum LARS algorithm for the Lasso path by quantising the search step of $i_{t+1}^{\rm join}$ and $\lambda_{t+1}^{\rm join}$ using the quantum minimum-finding subroutine from Dürr and Høyer [DH96] (Fact 10). For this, we input the vectors $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into KP-tree data structures accessed by QRAMs. Moreover, as described in Section 2.3, we assume the matrix $\mathbf{X}$ is stored in a QROM. The final result is an improvement in the runtime over the usual LARS algorithm to $\widetilde{O}(n\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2})$ per iteration.
1 Input: Vector $\mathbf{y}\in\mathbb{R}^{n}$, matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\delta\in(0,1)$, and $T\in\mathbb{N}$ Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0}$, $t=0$ while $\mathcal{I}\neq\varnothing$, $t\leq T$ do
       $\bm{\eta}_{\mathcal{A}}\leftarrow\operatorname{sign}\big{(}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\big{)}$ $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$   // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
       $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$   // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
2       Compute $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ classically and input them into QRAMs Define $\Lambda_{i}^{\rm join}\triangleq\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and $\Lambda_{i}^{\rm cross}\triangleq\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]$ $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ $i_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ and $\lambda_{t+1}^{\rm join}\leftarrow\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ with failure probability $\frac{\delta}{T}$ (Fact 10) $\lambda_{t+1}\leftarrow\max\{\lambda_{t+1}^{\rm join},\lambda_{t+1}^{\rm cross}\}$ $\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$ if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then
@@ 6      $t\leftarrow t+1$
Output: Coefficients $[(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\hat{\bm{\beta}}(\lambda_{1})),\dots]$
@@
$O(n\sqrt{|\mathcal{I}|}\log(T/\delta)\operatorname{poly}\log(nd)+n|\mathcal{A}|+|\mathcal{A}|^{2})$
@@
The time complexity of Algorithm 2 is basically the same as Algorithm 1, the only difference being that uploading $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into KP-trees takes $O(n\log{n})$ time, and finding $i_{t+1}^{\rm join}$ and $\lambda_{t+1}^{\rm join}$ now requires $O(n\sqrt{|\mathcal{I}|}\log(T/\delta)\operatorname{poly}\log(nd))$ time, since that accessing $\mathbf{X}_{i}$ and computing $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ requires $O(n\operatorname{poly}\log(nd))$ and $O(n)$ time, respectively. ∎
@@ 4.2 Approximate quantum LARS algorithm
We now improve our previous simple quantum LARS algorithm. The main idea is to approximately compute the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})$, see Eq. (12), within the quantum minimum-finding algorithm instead of exactly computing them. This will lead to a quadratic improvement on the $n$ dependence in the term $\widetilde{O}(n\sqrt{|\mathcal{I}|})$ to $\widetilde{O}(\sqrt{n|\mathcal{I}|})$. Such approximation, however, hinders our ability to exactly find the joining variables $i_{t+1}^{\rm join}$ and points $\lambda_{t+1}^{\rm join}$. We can now only obtain a joining point $\widetilde{\lambda}_{t+1}^{\rm join}$ which is $\epsilon$-close to the true joining point $\lambda_{t+1}^{\rm join}$. This imposes new complications on the correctness analysis of the LARS algorithm, since we can no longer guarantee that the KKT conditions are satisfied. In order to tackle this issue, we consider an approximate version of the KKT conditions. We note that a similar concept was already introduced in [MY12].
@@
[…]
@@
The primal problem $\min_{\bm{\beta}\in\mathbb{R}^{d},{\bf z}\in\mathbb{R}^{n}}\frac{1}{2}\|\mathbf{y}-{\bf z}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{s.t.}~{}{\bf z}=\mathbf{X}\bm{\beta}$, with dummy variable $\bf z$, leads to the dual problem
$\max_{\bm{\kappa}\in\mathbb{R}^{n}}-\frac{1}{2}\bm{\kappa}^{\top}\bm{\kappa}-\bm{\kappa}^{\top}\mathbf{y}\quad\text{s.t.}~{}\|\mathbf{X}^{\top}\bm{\kappa}\|_{\infty}\leq\lambda.$
@@
is called a duality gap and provides an optimality guarantee (see e.g. [BL06, Section 4.3]):
@@
$(\widetilde{\bm{\kappa}}+\mathbf{y})^{\top}\widetilde{\bm{\kappa}}=\widetilde{\bm{\beta}}^{\top}\mathbf{X}^{\top}(\mathbf{X}\widetilde{\bm{\beta}}-\mathbf{y})=\sum_{i=1}^{d}|\widetilde{\beta}_{i}|\mathbf{X}_{i}^{\top}(\mathbf{X}\widetilde{\bm{\beta}}-\mathbf{y})\operatorname{sign}(\widetilde{\beta}_{i})\leq-(\lambda-\lambda\epsilon)\|\widetilde{\bm{\beta}}\|_{1},$
@@
1 Input: Vector $\mathbf{y}\in\mathbb{R}^{n}$, matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\delta\in(0,1)$, $\epsilon\in(0,2)$, and $T\in\mathbb{N}$ Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\widetilde{\bm{\beta}}(\lambda_{0})=0$, $t=0$ while $\mathcal{I}\neq\varnothing$, $t\leq T$ do
       $\bm{\eta}_{\mathcal{A}}\leftarrow\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))$ $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$   // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
       $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$   // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
2       Classically compute $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ and input them into QRAMs and KP-trees Define $\Lambda_{i}^{\rm join}\triangleq\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and $\Lambda_{i}^{\rm cross}\triangleq\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]$ $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ Obtain $\widetilde{i}_{t+1}^{\rm join}\in\big{\{}j\in\mathcal{I}:\Lambda_{j}^{\rm join}\geq(1-\epsilon/2)\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}\big{\}}$ with failure probability $\frac{\delta}{T}$ (Fact 11 and Lemma 13) $\widetilde{\lambda}_{t+1}^{\rm join}\leftarrow(1-\epsilon/2)^{-1}\Lambda_{\widetilde{i}_{t+1}^{\rm join}}^{\rm join}$ $\lambda_{t+1}\leftarrow\min\{\lambda_{t},\max\{\lambda_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}\}$ $\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$ if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then

### v2

@@ 4 Quantum algorithms
In this section, we introduce our quantum algorithms based on the LARS algorithm. In Section 4.1, we propose our simple quantum LARS algorithm that exactly computes the pathwise Lasso, while in Section 4.2 we improve upon this simple algorithm and propose the approximate quantum LARS algorithm. In Section 4.3, we dequantise the approximate LARS algorithm using sampling techniques. Finally, in Section 4.4, we analyse the algorithms’ complexity for the case when $\mathbf{X}$ is a random matrix from the $\mathbf{\Sigma}$-Gaussian ensemble.
@@ 4.1 Simple quantum LARS algorithm
[…]
1
Input: $T\in\mathbb{N}$, $\delta\in(0,1)$, $\mathbf{y}\in\mathbb{R}^{n}$, and $\mathbf{X}\in\mathbb{R}^{n\times d}$
2
Output: Exact regularisation path $\mathcal{P}$ with $T$ kinks with probability $\geq 1-\delta$
3
4Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0}$, $t=0$
5while $\mathcal{I}\neq\varnothing$, $t\leq T$ do
6       $\bm{\eta}_{\mathcal{A}}\leftarrow\operatorname{sign}\big{(}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\big{)}$
      $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$
        // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
7
      $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$
        // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
8
9      Compute $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ classically and input them into QRAMs
10      Define $\Lambda_{i}^{\rm join}\triangleq\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and $\Lambda_{i}^{\rm cross}\triangleq\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]$
11      $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$
12      $i_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ and $\lambda_{t+1}^{\rm join}\leftarrow\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}$ with failure probability $\frac{\delta}{T}$ (Fact 13)
13      $\lambda_{t+1}\leftarrow\max\{\lambda_{t+1}^{\rm join},\lambda_{t+1}^{\rm cross}\}$
14
15      $\hat{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$
16
17      if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then
18
@@ 24      $t\leftarrow t+1$
return coefficients $[(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})),(\lambda_{1},\hat{\bm{\beta}}(\lambda_{1})),\dots]$
@@
$O\big{(}n\sqrt{|\mathcal{I}|}\log(T/\delta)\operatorname{poly}\log(nd)+n|\mathcal{A}|+|\mathcal{A}|^{2}\big{)}$
@@
The time complexity of Algorithm 2 is basically the same as Algorithm 1, the only difference being that uploading $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into KP-trees takes $O(n\log{n})$ time, and finding $i_{t+1}^{\rm join}$ and $\lambda_{t+1}^{\rm join}$ now requires $O\big{(}n\sqrt{|\mathcal{I}|}\log(T/\delta)\operatorname{poly}\log(nd)\big{)}$ time, since that accessing $\mathbf{X}_{i}$ and computing $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ requires $O(n\operatorname{poly}\log(nd))$ and $O(n)$ time, respectively. ∎
@@ 4.2 Approximate quantum LARS algorithm
[…]
@@
$\lambda(1-\epsilon)\leq\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}})\operatorname{sign}(\widetilde{\beta}_{j})\leq\lambda$ $\quad\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \widetilde{\beta}_{j}\neq 0,$
$|\mathbf{X}_{j}^{\top}(\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}})|\leq\lambda$ $\quad\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ \widetilde{\beta}_{j}=0.$
@@
The primal problem $\min_{\bm{\beta}\in\mathbb{R}^{d},{\bf z}\in\mathbb{R}^{n}}\frac{1}{2}\|\mathbf{y}-{\bf z}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\leavevmode\nobreak\ \text{s.t.}\leavevmode\nobreak\ {\bf z}=\mathbf{X}\bm{\beta}$, with dummy variable $\bf z$, leads to the dual problem
$\max_{\bm{\kappa}\in\mathbb{R}^{n}}-\frac{1}{2}\bm{\kappa}^{\top}\bm{\kappa}-\bm{\kappa}^{\top}\mathbf{y}\quad\text{s.t.}\leavevmode\nobreak\ \|\mathbf{X}^{\top}\bm{\kappa}\|_{\infty}\leq\lambda.$
@@
is called a duality gap and provides an optimality guarantee (see, e.g., [12, Section 4.3]):
@@
$(\widetilde{\bm{\kappa}}+\mathbf{y})^{\top}\widetilde{\bm{\kappa}}=\widetilde{\bm{\beta}}^{\top}\mathbf{X}^{\top}(\mathbf{X}\widetilde{\bm{\beta}}-\mathbf{y})=\sum_{i=1}^{d}|\widetilde{\beta}_{i}|\mathbf{X}_{i}^{\top}(\mathbf{X}\widetilde{\bm{\beta}}-\mathbf{y})\operatorname{sign}(\widetilde{\beta}_{i})\leq-\lambda(1-\epsilon)\|\widetilde{\bm{\beta}}\|_{1},$
@@
1
Input: $T\in\mathbb{N}$, $\delta\in(0,1)$, $\epsilon>0$, $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times d}$, and $\{\alpha_{\mathcal{A}}\geq\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\}_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}$
2
Output: Regularisation path $\widetilde{\mathcal{P}}$ with $T$ kinks and error $\lambda\epsilon\|\widetilde{\bm{\beta}}(\lambda)\|_{1}$ with probability $\geq 1-\delta$
3
4Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\widetilde{\bm{\beta}}(\lambda_{0})=0$, $t=0$
5while $\mathcal{I}\neq\varnothing$, $t\leq T$ do
6       $\bm{\eta}_{\mathcal{A}}\leftarrow\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))$
      $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$
        // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
7
      $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$
        // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
8
9      Classically compute $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ and input them into QRAMs and KP-trees
10      Define $\Lambda_{i}^{\rm join}\triangleq\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and $\Lambda_{i}^{\rm cross}\triangleq\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]$
11      $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$
12      Obtain $\widetilde{i}_{t+1}^{\rm join}\in\big{\{}j\in\mathcal{I}:\Lambda_{j}^{\rm join}\geq\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}\}\big{\}}$ with failure probability $\frac{\delta}{T}$ (Fact 14 and Lemma 16)
13      $\widetilde{\lambda}_{t+1}^{\rm join}\leftarrow\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}^{-1}\Lambda_{\widetilde{i}_{t+1}^{\rm join}}^{\rm join}$
14
15      $\lambda_{t+1}\leftarrow\min\{\lambda_{t},\max\{\lambda_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}\}$
16
17      $\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$
18
19      if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then